cyclist travels from town A to town C through another town B. He travels A to B at a speed of 6 km/h and B to C at a speed of 9 km/h. The time taken for journey is 4 hrs. On his return journey, the travels at a speed of 6 km from C to B and at a speed of 9 km/h. The time taken for return journey is 4 hrs and 20 minutes. Find the distance AC and BC.
1
Expert's answer
2020-08-10T17:43:58-0400

Let the distance from A to B =X km and from B to C =Y km.

On the going journey ;Time taken from A to B="\frac{X} {6}" hrs and B to C="\frac{Y} {9}" hrs.

Hence ; "\frac{x} {6}" +"\frac{y} {9}" = 4hrs ............ (i)

On the Return journey ;Time taken from C to B ="\frac{y} {6}" hrs and B to A ="\frac{y} {9}" hrs.

Hence ;"\frac{y} {6}" +"\frac{y} {9}" ="\frac{13} {3}" hrs........... (ii)


Making y the subject of the formula in equation (i) ;

y =(4-"\frac{x} {6}") 9=

y=36-"\frac{3} {2}" x

Substituting y in the second equation ;

"(\frac{36-1.5x}{6})" +"\frac{x} {9}" ="\frac{13}{3}"


=108-4.5x+2x=78

=-2.5x=-30

x=12

y=36-(1.5×12)=18

Hence Distance AC=12+18=30km

Distance BC =18km